A cylindrical tank has a height of 6.0 ft and a diameter of 3.0 ft. What is its capacity in liters if the volume of a cylinder can be found by the equation 2π r 2 h? Use: 3.28 ft = 1.0 m; 1000 L = 1 m3
1
Expert's answer
2020-10-02T07:14:13-0400

"6ft=1.8288m"


"3ft=0.9144m"


"V=\pi r^2h=3.14\cdot(0.9144/2)^2\cdot1.8288\approx1.2(m^3)=1200(L)"